OpenStudy (anonymous):

3√45x^3 + 2√12 + √27 - 3√20x^3

OpenStudy (anonymous):

\[9\sqrt{5x^3} + 4\sqrt{3} + 3\sqrt{3} - 6\sqrt{5x^3} = 7\sqrt{3} + 3\sqrt{5x^3}\]

OpenStudy (anonymous):

When you siimplify the radical for instance \[2\sqrt{12}\] first rewrite 12 as \[2 * \sqrt{4 * 3} = 2*\sqrt{4}*\sqrt{3} = 2*2*\sqrt{3} = 4\sqrt{3}\] Thats how you simplify the radicals when you have a number in front

OpenStudy (anonymous):

Thank you. But the answers I can choose from are: a. 40x√x b. 3x√5x + 7√3 c. 3x√10x + 7√6 d. 10x√8x

OpenStudy (anonymous):

2√12 = 2*2*√3 = 4√3 √27 = 3√3 3√45x^3 = 9√5 x^3 -3√20x^3 = - 6√5x^3 adding you get 3√5x^3 + 7√3

OpenStudy (anonymous):

oh its b then the x^3 is contained in the square root